OpenStudy (anonymous):

help! finding a taylor polynomial of f(x,y)= sin(x+y)/x+y at point (0,0)

OpenStudy (amistre64):

hmm, i ve never tried to do a multivariable taylor before; im sure its based on teh same principles tho

OpenStudy (amistre64):

gonna have to download the chrome to be of any use on this site ... i need to read up on this a bit as well

OpenStudy (anonymous):

interesting thats awesum enough

OpenStudy (amistre64):

f fx fxx fy fyy fxy are all the taylor stuff that correspond to their variable parts too

OpenStudy (amistre64):

\[ f= sin(x+y)\ (x+y)^{-1}\] \[ fx= cos(x+y)\ (x+y)^{-1}+sin(x+y)\ -(x+y)^{-2}\] i have got to write this down on paper; cant keep track of it in this accursed latex coding

OpenStudy (amistre64):

\[f= \frac{sin(x+y)}{(x+y)}\] \[fx= \frac{cos(x+y)}{(x+y)}-\frac{sin(x+y)}{(x+y)^2}\] \[fxx= -\frac{sin(x+y)}{(x+y)}-2\frac{cos(x+y)}{(x+y)^2}+2\frac{sin(x+y)}{(x+y)^3}\] now for the y parts

OpenStudy (amistre64):

actually, the y parts i beleive are the exact same things as the x parts since there is no variation among the variables that would consititute a major change up. so on to the xy part :)

OpenStudy (amistre64):

yep, since all we do is pop out +1s all over the place the xy parts the same as an xx part as well

OpenStudy (amistre64):

so we double up the fx: fx fy we get to more versions of fxx: fxx fyy fxy and we should be good to go

OpenStudy (amistre64):

the issue of course now is that there is no place for 0,0 to be in the denominator since /0 is bad mojo

OpenStudy (anonymous):

ow..thats interesting

OpenStudy (amistre64):

i checked with the wolf and it likes my deriving :) but the issue I still have is in determining values using the point 0,0

OpenStudy (anonymous):

if its a third degree what happens?

OpenStudy (amistre64):

we still get the x+y on the bottom so its really not gonna matter how many iterations we do; they will all fail once you put in the 0,0

OpenStudy (amistre64):

http://web.monroecc.edu/manila/webfiles/calcNSF/JavaCode/CalcPlot3D.htm we might be able to take the limit as this apporaches 0,0

OpenStudy (amistre64):

in that sense it approaches 1

OpenStudy (anonymous):

lim (x,y) -(pi/2;pi/2) of cosxsiny+ytanx

OpenStudy (amistre64):

cos(pi/2) sin(pi/2) = 0 so thats not an issue

OpenStudy (amistre64):

y going to pi/2 isnt an issue either but might play into the picture if tanx dont go bad

OpenStudy (amistre64):

sin(x) ----- as x goes to pi/2 hmmm cos(x)

OpenStudy (amistre64):

i beleive that blows up into infinity but id have to chk with the wolf to be certain

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=limit+tan%28x%29+x+to+pi%2F2 yeah, so the limit does not exist

OpenStudy (anonymous):

thank you that helped